# Accelerating Image Generation with Sub-path Linear Approximation Model

### v2

@@ Abstract
Diffusion models have significantly advanced the state of the art in image, audio, and video generation tasks. However, their applications in practical scenarios are hindered by slow inference speed. Drawing inspiration from the approximation strategies utilized in consistency models, we propose the Sub-path Linear Approximation Model (SLAM), which accelerates diffusion models while maintaining high-quality image generation. SLAM treats the PF-ODE trajectory as a series of PF-ODE sub-paths divided by sampled points, and harnesses sub-path linear (SL) ODEs to form a progressive and continuous error estimation along each individual PF-ODE sub-path. The optimization on such SL-ODEs allows SLAM to construct denoising mappings with smaller cumulative approximated errors. An efficient distillation method is also developed to facilitate the incorporation of more advanced diffusion models, such as latent diffusion models. Our extensive experimental results demonstrate that SLAM achieves an efficient training regimen, requiring only 6 A100 GPU days to produce a high-quality generative model capable of 2 to 4-step generation with high performance. Comprehensive evaluations on LAION, MS COCO 2014, and MS COCO 2017 datasets also illustrate that SLAM surpasses existing acceleration methods in few-step generation tasks, achieving state-of-the-art performance both on FID and the quality of the generated images. Project Page: https://subpath-linear-approx-model.github.io/. .
@@ 1 Introduction
Diffusion models, also known as score-based generative models, have emerged as a potent paradigm in generative computer vision, enabling the synthesis of highly realistic images by progressively refining random noise into structured visual content [9, 46, 29, 32, 45]. Despite their impressive ability, one of the primary challenges associated with diffusion models lies in their computational intensity, often requiring hundreds of iterative steps to produce a single image. This has spurred a surge of research focused on accelerating diffusion models to retain high-quality outputs while significantly reducing the computation cost during the inference phase [42, 23, 24, 36, 50, 49, 21, 44, 26, 20].
@@
To address these challenges, we propose a novel approach in this paper, designated as the Sub-path Linear Approximation Model (SLAM). SLAM adheres to the foundational concept of cumulative approximation of PF-ODE trajectories but innovates through its sustained learning from Sub-path Linear (SL) ODEs. Specifically, we dissect the sub-path learning objective based on the noise prediction design [9, 14] into two interrelated aspects, and establish the SL-ODEs to give respective progressive or continuous estimation for each component, by a carefully designed linear interpolation between the endpoints of sub-paths. We then utilize the SL-ODEs to approximate the complete PF-ODE trajectories which allows a more nuanced optimization. Consequently, the prediction error of our approach is assessed through iterative solutions of all SL-ODEs, enabling a reduction of cumulative errors and an enhancement in image generation quality. Furthermore, we also develop an efficient distillation procedure for our SLAM which enables the incorporation with pre-trained latent diffusion models [34] (e.g., Stable Diffusion). Our contributions can be summarized as below:
@@
    To address these challenges, we propose a novel approach as Sub-path Linear Approximation Model (SLAM). SLAM employs Sub-path Linear (SL) ODEs to continuously approximate the complete PF-ODE trajectories and progressively optimize the sub-path learning objectives through the SL-ODE solutions, which could construct the denoising mappings with smaller cumulative approximated errors.
@@
    Leveraging the proposed SLAM and SL-ODE framework, we put forth an efficient distillation method to train the model using pre-trained latent diffusion models. When integrated with powerful pre-trained models like Stable Diffusion, our approach allows a more efficient training and respectively attains impressive FIDs as 10.09, 10.06, 20.77 in LAION, MS COCO 2014, MS COCO 2017 datasets, achieving better performance and close inference latency to all previous accelerating approaches.
Figure 1: Our Sub-path Linear Approximation Model employs Sub-path Linear ODEs to approximate the sub-paths on the PF-ODE trajectories, which is determined by the linear interpolation of corresponding endpoints. SLAM is then trained based on the consistent mapping along SL-ODEs to minimize the approximated errors.
@@ 3 Preliminaries
Diffusion Based Models are a class of generative models that gradually transform data into a noisy state through Gaussian perturbations and subsequently learn to reverse this process to reconstruct the original data by progressively denoising it. Denote $\bm{x}_{0}$ as the data sampled from the original distribution $\bm{x}_{0}\sim p_{data}(\bm{x})$ and $\alpha(t),\sigma(t)$ as functions that define a noise schedule. Diffusion models transition the data to a noise-corrupted marginal distribution, which can be expressed as:
@@ 4.1 Approximation Strategy for Denoiser
One-step Denoiser Parameterization. To synthesize an image from a sampled input $\bm{x}_{t}$ at a large time step $t$ in one-step, a natural approach is to adopt the strategy from [9] that employs a neural network $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$ to predict a standard Gaussian distribution, which implements the denoising mapping parameterized as $\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)=\frac{\bm{x}_{t}-\sigma(t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)}{\alpha(t)}$. By redefining the target distribution for $(\bm{x}_{t},t)$ as $\bm{x}^{t}_{0}=\alpha(t)\bm{x}_{0}\sim p_{data,t}(\alpha(t)\bm{x})$ and $D_{\bm{\theta}}(\bm{x}_{t},t)=\alpha(t)*\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)=\bm{x}_{t}-\sigma(t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$, this predictive formulation can be recast into a canonical denoiser function defined in [14] that aims to minimize the denoising error as follows:
$\mathcal{L}_{D}({\bm{\theta}})=\mathbb{E}_{\bm{x}^{t}_{0}\sim p_{data,t},\bm{x}_{t}\sim\mathcal{N}(\bm{x}^{t}_{0},\sigma(t)^{2}I)}[|D_{\bm{\theta}}(\bm{x}_{t},t)-\alpha(t)\bm{x}_{0}|],$ (8)
@@
$\mathbb{E}[|D_{\bm{\theta}}(\bm{x}_{t},t)-\alpha(t)\bm{f}_{\bm{\theta}}(\bm{x}_{t-k},t-k)|].$ (9)
By incorporating the expressions for $\bm{f}_{\bm{\theta}}(\bm{x}_{t-k},t-k)$ and $D_{\bm{\theta}}(\bm{x}_{t},t)$, we derive the approximated error estimation based on $\bm{\epsilon}_{\bm{\theta}}(\cdot,\cdot)$ as:
@@
where the mentioned impact on optimization is reduced as the coefficient is amplified by $\alpha(t-k)$. When $k$ is limited to 1, the error between the mapping result $\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)$ and the trajectory origin $\bm{x}_{0}$ can be quantified by the accumulation of incremental approximated errors [44]: $|\bm{x}_{0}-\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)|\leq\sum\limits_{1\leq t^{\prime}\leq t}|\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}},t^{\prime})-\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}-1},t^{\prime}-1)|$. Ideally if the error of one single approximation can be bounded, we can reduce the cumulative error by decreasing the number of approximations. This technique, also called Skipping-Step in LCM [26], extends to optimize the error for skipping sampled points on the trajectories as $|\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}},t^{\prime})-\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}-k},t^{\prime}-k)|$ for a fixed skipping step size $k$. However, our insights reveal this precondition does not hold for extended situations. Denote $\{\bm{x}_{t^{\prime}}\}_{t^{\prime}\in[t-k,t]}$ as the sub-path between $\bm{x}_{t-k}$ and $\bm{x}_{t}$ from the original PF-ODE trajectory, we discern that the learning objective in Eq. 10 for $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$ can be decomposed into two complementary components: 1) $dist_{\Delta}(\bm{x}_{t},\bm{x}_{t-k},t)=\bm{x}_{t}-\frac{\alpha(t)}{\alpha(t-k)}\bm{x}_{t-k}$, which gauges the incremental distance from $\bm{x}_{t-k}$ to $\bm{x}_{t}$ attributable to the drift and diffusion processes, and 2) $dist_{0,{\bm{\theta}}}(\bm{x}_{t-k},t-k,t)=\frac{\alpha(t)}{\alpha(t-k)}\sigma(t-k)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t-k},t-k)$, which captures the denoising contribution from previous time steps that should be coherently propagated to subsequent time steps $t$. Thus we rewrite Eq. 10 as a sub-path learning objective:
@@ 4.2 Sub-path Linear Approximation Model
To improve the learning objective in Eq. 11, in this paper we introduce a new approach for accelerating diffusion models termed Sub-path Linear Approximation Model (SLAM). SLAM introduces Sub-path Linear (SL) ODEs to approximate the sub-paths on the PF-ODE trajectories as a linear interpolation between the according sub-path endpoints. As the optimization based on such SL-ODEs gives a respectively progressive and continuous estimation for the decomposed two terms in Eq. 11, our SLAM is trained based on the conducted SL-ODE learning objectives, and achieves smaller overall prediction errors and better generation quality. We also develop an efficient distillation procedure for latent diffusion models  [34], with Multiple Estimation strategy which improves the estimated results of teacher models.
@@ 4.2.1 Sub-path Linear ODE
$\bm{x}_{\gamma,t}$ $=\frac{\alpha(t)}{\alpha(t-k)}\bm{x}_{t-k}+\gamma*dist_{\Delta}(\bm{x}_{t},\bm{x}_{t-k},t),$ (12)
@@
$D_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t)=\bm{x}_{\gamma,t}-\sigma(\gamma,t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t),$ (14)
$\bm{f}_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t)=\frac{D_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t)}{\alpha(t)}.$
Incorporating these into Eq. 11, we derive the sub-path learning objective for our SLAM model as :
$\small\mathcal{L}_{\text{SLAM}}({\bm{\theta}},k)=\mathbb{E}[|\gamma*dist_{\Delta}(\bm{x}_{t},\bm{x}_{t-k},t)+dist_{0,{\bm{\theta}}}(\bm{x}_{t-k},t,t-k)-\sigma(r,t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{\gamma,t},t)|],$ (15)
which gives a progressive estimation for the otherwise intractable $dist_{\Delta}$ objective. The value for $\sigma(\gamma,t)$ can be precisely estimated from the distribution $p_{t}(\bm{x}_{t})$ and $p_{t-k}(\bm{x}_{t-k})$ but has a complex expression. Empirically we utilize an approximate result as $\sigma(\gamma,t)=(1-\gamma)\frac{\alpha(t)}{\alpha(t-k)}\sigma(t-k)+\gamma*\sigma(t)$. Compared to consistency models which adopt Eq. 10 or Eq. 11, our $\mathcal{L}_{\text{SLAM}}$ maintains a progressive estimation for $dist_{\Delta}$ and a consistent estimation for $dist_{0,{\bm{\theta}}}$, which enables the learning for large skipping step size. The overall prediction error can still be assessed by the aggregate of approximated errors between the sub-path endpoints and the approximated error between these points is continuously optimized through the SL-ODEs. Consequently, the optimization for the approximated errors in our SLAM could be significantly improved. Our approach could further benefit from the increased skipping step size, allowing our method to generate images of higher quality with reduced sampling steps in a more efficient training.
@@ 4.2.2 Sub-path Linear Approximation Distillation
In this paper, we adopt pre-trained Stable Diffusion (SD) models [34] to obtain the solution PF-ODE trajectories which we build our SL-ODEs upon, and we call the approach Sub-path Linear Approximation Distillation (SLAD). To achieve conditional generation with the conditional input $c$, the noise prediction model is parameterized as $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},c,t)$ [46, 23]. We also introduce $\gamma$ into the prediction models for solving our SL-ODEs, and leverage the $\gamma-$conditioned training where $\gamma$ is converted to Fourier embeddings and fed into the models as an input. Specifically, to predict $\bm{z}_{0}$ in the latent space, the generation function for SLAM is defined as:
@@
where $\bm{f}_{\bm{\theta}}(\bm{z}_{\gamma,t},c,\gamma,t)$ mirrors Eq. 14 while replacing $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{\gamma,t},\gamma,t)$ with the conditional form $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{\gamma,t},c,\gamma,t)$. The functions $c_{\text{skip}}$ and $c_{\text{out}}$ are leveraged to ensure that $\bm{F}_{\bm{\theta}}(\bm{z}_{1,0},c,1,0)\equiv\bm{z}_{0}$ (we regard $F_{\bm{\theta}}$ as the same expression of $f_{\bm{\theta}}$ since $c_{\text{skip}}(t)\ll c_{\text{out}}(t)$ for most time steps). Integrating this with Eq. 9, our SLAD approach minimizes the following objective:
$\small\mathcal{L}_{\text{SLAD}}({\bm{\theta}},{\bm{\theta}}^{-};\phi)=\mathbb{E}_{\bm{z}_{0}\sim p_{data},t\sim\mathcal{U}[k,T],\gamma\sim\mathcal{U}[0,1]}[|\bm{F}_{\bm{\theta}}(\bm{z}_{\gamma,t},c,\gamma,t)-\bm{F}_{{\bm{\theta}}^{-}}(\hat{\bm{z}}^{\Phi}_{1,t-k},c,1,t-k)|],$ (17)
@@
Algorithm 1 Sub-path Linear Approximation Distillation (SLAD)
@@
The pseudo-code for SLAD is presented in Algorithm 1. SLAD shares a similar training pipeline with consistency models[44, 26] but can be distinguished as it optimizes the sub-path learning objectives based on the SL-ODEs and utilizes the $\gamma$-conditioned training. For a pair of input noise and time step $(\bm{z}_{t},t)$, SLAM gives the prediction of the denoised latent $\hat{\bm{z}}_{0}$ as:
@@
for one-step generation, adhering strictly to the $\gamma=1$ condition. We also use the same iterative sample strategy as illustrated in [44] which could improve the quality of the generated images. In practice, we set the $-\gamma\text{-}$embedding to $\mathbf{0}$ for $\gamma=1$, thereby allowing the weights associated with trained $\gamma$-embeddings to be discarded post-training. Thus our Sub-path Linear Approximation Model (SLAM) necessitates no additional parameters beyond the training phase and can be utilized the same as teacher models.

### v3

# https://subpath-linear-approx-model.github.io/

@@ Abstract
Diffusion models have significantly advanced the state of the art in image, audio, and video generation tasks. However, their applications in practical scenarios are hindered by slow inference speed. Drawing inspiration from the consistency models, we propose the Sub-Path Linear Approximation Model (SPLAM), which can accelerate diffusion models while maintaining high-quality image generation. SPLAM treats the PF-ODE trajectory as a series of PF-ODE sub-paths divided by sampled points, and harnesses sub-path linear (SL) ODEs to form a progressive and continuous error estimation along each individual PF-ODE sub-path. The optimization on such SL-ODEs allows SPLAM to construct denoising mapping with smaller cumulative approximated error. An efficient distillation method is also developed to facilitate the incorporation of pre-trained diffusion models, such as latent diffusion models. The extensive experimental results demonstrate SPLAM achieves remarkable training efficiency, requiring only 6 A100 GPU days to produce a high-quality generative model capable of 2 to 4-step generation. Comprehensive evaluations on LAION, MS COCO 2014, and MS COCO 2017 datasets also illustrate that SPLAM surpasses the existing acceleration methods in few-step generation tasks, achieving state-of-the-art performance both on FID and the quality of the generated images. .
@@ 1 Introduction
Diffusion models, also known as score-based generative models, have emerged as a potent paradigm in generative computer vision, enabling the synthesis of highly realistic images by progressively refining random noise into structured visual content [9, 46, 29, 32, 45]. Despite their impressive ability, one of the primary challenges associated with diffusion models lies in their computational intensity, often requiring hundreds of iteration steps to produce a single image. This has spurred a surge of research focused on accelerating diffusion models to retain high-quality outputs while significantly reducing the computation cost during the inference phase [42, 23, 24, 36, 50, 49, 21, 44, 26, 20].
@@
To address these challenges, we propose a novel approach in this paper, designated as the Sub-Path Linear Approximation Model (SPLAM). SPLAM adheres to the foundational concept of cumulative approximation of PF-ODE trajectories but innovates through its sustained learning from Sub-Path Linear (SL) ODEs. Specifically, we dissect the sub-path learning objective based on the noise prediction design [9, 14] into two interrelated aspects, and establish the SL-ODEs to give respective progressive or continuous estimation for each component, by a carefully designed linear interpolation between the endpoints of sub-paths. We then utilize the SL-ODEs to approximate the complete PF-ODE trajectories which allows a more nuanced optimization. Consequently, the prediction error of our approach is assessed through iterative solutions of all SL-ODEs, enabling a reduction of cumulative errors and an enhancement in image generation quality. Furthermore, we also develop an efficient distillation procedure for our SPLAM which enables the incorporation with pre-trained latent diffusion models [34] (e.g., Stable Diffusion). Our contributions can be summarized as below:
@@
    To address these challenges, we propose a novel approach as Sub-Path Linear Approximation Model (SPLAM). SPLAM employs Sub-Path Linear (SL) ODEs to continuously approximate the complete PF-ODE trajectories and progressively optimize the sub-path learning objectives, which could construct the denoising mappings with smaller cumulative approximated errors.
@@
    Leveraging the proposed SPLAM and SL-ODE framework, we put forth an efficient distillation method. When integrated with powerful pre-trained models like Stable Diffusion, our approach allows more efficient training and respectively attains impressive FIDs as 10.09, 10.06, 20.77 in LAION, MS COCO 2014, MS COCO 2017 datasets, achieving better performance and close inference latency to all previous accelerating approaches.
Figure 1: Our Sub-Path Linear Approximation Model employs Sub-Path Linear ODEs to approximate the sub-paths on the PF-ODE trajectories, which is determined by the linear interpolation of corresponding endpoints. SPLAM is then trained based on the consistent mapping along SL-ODEs to minimize the approximated errors.
@@ 3 Preliminaries
Diffusion Models are a class of generative models that gradually transform data into a noisy state through Gaussian perturbations and subsequently learn to reverse this process to reconstruct the original data by progressively denoising it. Denote $\bm{x}_{0}$ as the data sampled from the original distribution $\bm{x}_{0}\sim p_{data}(\bm{x})$ and $\alpha(t),\sigma(t)$ as functions that define a noise schedule. Diffusion models transition the data to a noise-corrupted marginal distribution, which can be expressed as:
@@ 4.1 Approximation Strategy for Denoiser
One-step Denoiser Parameterization. To synthesize an image from a sampled input $\bm{x}_{t}$ at a large time step $t$ in one-step, a natural approach is to adopt the strategy from [9] that employs a neural network $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$ to predict a standard Gaussian distribution, which implements the denoising mapping parameterized as $\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)=\frac{\bm{x}_{t}-\sigma(t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)}{\alpha(t)}$. By redefining the target distribution for $(\bm{x}_{t},t)$ as $\bm{x}^{t}_{0}=\alpha(t)\bm{x}_{0}\sim p_{data,t}(\alpha(t)\bm{x})$ and $\bm{D}_{\bm{\theta}}(\bm{x}_{t},t)=\alpha(t)*\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)=\bm{x}_{t}-\sigma(t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$, this predictive formulation can be recast into a canonical denoiser function defined in [14] that aims to minimize the denoising error as follows:
$\mathcal{L}_{\bm{D}}({\bm{\theta}})=\mathbb{E}_{\bm{x}^{t}_{0}\sim p_{data,t},\bm{x}_{t}\sim\mathcal{N}(\bm{x}^{t}_{0},\sigma(t)^{2}I)}[|\bm{D}_{\bm{\theta}}(\bm{x}_{t},t)-\alpha(t)\bm{x}_{0}|],$ (8)
@@
$\mathbb{E}[|\bm{D}_{\bm{\theta}}(\bm{x}_{t},t)-\alpha(t)\bm{f}_{\bm{\theta}}(\bm{x}_{t-k},t-k)|].$ (9)
By incorporating the expressions for $\bm{f}_{\bm{\theta}}(\bm{x}_{t-k},t-k)$ and $\bm{D}_{\bm{\theta}}(\bm{x}_{t},t)$, we derive the approximated error estimation based on $\bm{\epsilon}_{\bm{\theta}}(\cdot,\cdot)$ as:
@@
where the mentioned impact on optimization is reduced as the coefficient is amplified by $\alpha(t-k)$. When $k$ is limited to 1, the error between the mapping result $\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)$ and the trajectory origin $\bm{x}_{0}$ can be quantified by the accumulation of incremental approximated errors [44]: $|\bm{x}_{0}-\bm{f}_{\bm{\theta}}(\bm{x}_{t},t)|\leq\sum\limits_{1\leq t^{\prime}\leq t}|\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}},t^{\prime})-\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}-1},t^{\prime}-1)|$. Ideally if the error of one single approximation can be bounded, we can reduce the cumulative error by decreasing the number of approximations. This technique, also called Skipping-Step in LCM [26], extends to optimize the error for skipping sampled points on the trajectories as $|\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}},t^{\prime})-\bm{f}_{\bm{\theta}}(\bm{x}_{t^{\prime}-k},t^{\prime}-k)|$ for a fixed skipping step size $k$. However, our insights reveal this precondition does not hold for extended situations. Denote $\{\bm{x}_{t^{\prime}}\}_{t^{\prime}\in[t-k,t]}$ as the sub-path between $\bm{x}_{t-k}$ and $\bm{x}_{t}$ from the original PF-ODE trajectory, we discern that the learning objective in Eq. 10 for $\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)$ can be decomposed into two complementary components: 1) $dist_{\Delta}(\bm{x}_{t-k},\bm{x}_{t},t)=\bm{x}_{t}-\frac{\alpha(t)}{\alpha(t-k)}\bm{x}_{t-k}$, which gauges the incremental distance from $\bm{x}_{t-k}$ to $\bm{x}_{t}$ attributable to the drift and diffusion processes, and 2) $dist_{0,{\bm{\theta}}}(\bm{x}_{t-k},t-k,t)=\frac{\alpha(t)}{\alpha(t-k)}\sigma(t-k)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t-k},t-k)$, which captures the denoising contribution from previous time steps that should be coherently propagated to subsequent time steps $t$. Thus we rewrite Eq. 10 as a sub-path learning objective:
@@ 4.2 Sub-Path Linear Approximation Model
To improve the learning objective in Eq. 11, in this paper we introduce a new approach for accelerating diffusion models termed Sub-path Linear Approximation Model (SPLAM). SPLAM introduces Sub-Path Linear (SL) ODEs to approximate the sub-paths on the PF-ODE trajectories as a linear interpolation between the according sub-path endpoints. As the optimization based on such SL-ODEs gives a respectively progressive and continuous estimation for the decomposed two terms in Eq. 11, our SPLAM is trained based on the conducted SL-ODE learning objectives, and achieves smaller overall prediction errors and better generation quality. We also develop an efficient distillation procedure for latent diffusion models  [34], with Multiple Estimation strategy which improves the estimated results of teacher models.
@@ 4.2.1 Sub-Path Linear ODE
$\bm{x}_{\gamma,t}$ $=\frac{\alpha(t)}{\alpha(t-k)}\bm{x}_{t-k}+\gamma*dist_{\Delta}(\bm{x}_{t},\bm{x}_{t-k},t)$ (12)
@@
$\bm{D}_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t)=\bm{x}_{\gamma,t}-\sigma(\gamma,t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t),$ (14)
$\bm{f}_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t)=\frac{\bm{D}_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t)}{\alpha(t)}.$
Incorporating these into Eq. 11, we derive the sub-path learning objective for our SPLAM model as :
$\small\mathcal{L}_{\text{SPLAM}}({\bm{\theta}},k)=\mathbb{E}[|\gamma*dist_{\Delta}(\bm{x}_{t},\bm{x}_{t-k},t)+dist_{0,{\bm{\theta}}}(\bm{x}_{t-k},t,t-k)-\sigma(\gamma,t)\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{\gamma,t},\gamma,t)|],$ (15)
which gives a progressive estimation for the otherwise intractable $dist_{\Delta}$ objective. The value for $\sigma(\gamma,t)$ can be precisely estimated from the distribution $p_{t}(\bm{x}_{t})$ and $p_{t-k}(\bm{x}_{t-k})$ but has a complex expression. Empirically we utilize an approximate result as $\sigma(\gamma,t)=(1-\gamma)\frac{\alpha(t)}{\alpha(t-k)}\sigma(t-k)+\gamma*\sigma(t)$. Compared to consistency models which adopt Eq. 10 or Eq. 11, our $\mathcal{L}_{\text{}}$ maintains a progressive estimation for $dist_{\Delta}$ and a consistent estimation for $dist_{0,{\bm{\theta}}}$, which enables the learning for large skipping step size. The overall prediction error can still be assessed by the aggregate of approximated errors between the sub-path endpoints and the approximated error between these points is continuously optimized through the SL-ODEs. Consequently, the optimization for the approximated errors in our SPLAM could be significantly improved. Our approach could further benefit from the increased skipping step size, allowing our method to generate images of higher quality with reduced sampling steps in more efficient training.
@@ 4.2.2 Sub-Path Linear Approximation Distillation
In this paper, we adopt pre-trained Stable Diffusion (SD) models [34] to obtain the solution PF-ODE trajectories which we build our SL-ODEs upon, and we call the approach Sub-path Linear Approximation Distillation (SPLAD). To achieve conditional generation with the conditional input $c$, the noise prediction model is parameterized as $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},c,t)$ [46, 23]. We also introduce $\gamma$ into the prediction models for solving our SL-ODEs, and leverage the $\gamma-$conditioned training where $\gamma$ is converted to Fourier embeddings and fed into the models as an input. Specifically, to predict $\bm{z}_{0}$ in the latent space, the generation function for SPLAM is defined as:
@@
where $\bm{f}_{\bm{\theta}}(\bm{z}_{\gamma,t},c,\gamma,t)$ mirrors Eq. 14 while replacing $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{\gamma,t},\gamma,t)$ with the conditional form $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{\gamma,t},c,\gamma,t)$. The functions $c_{\text{skip}}$ and $c_{\text{out}}$ are leveraged to ensure that $\bm{F}_{\bm{\theta}}(\bm{z}_{1,0},c,1,0)\equiv\bm{z}_{0}$ (we regard $F_{\bm{\theta}}$ as the same expression of $f_{\bm{\theta}}$ since $c_{\text{skip}}(t)\ll c_{\text{out}}(t)$ for most time steps). Integrating this with Eq. 9, our SPLAD approach minimizes the following objective:
$\small\mathcal{L}_{\text{SPLAD}}({\bm{\theta}},{\bm{\theta}}^{-};\phi)=\mathbb{E}_{\bm{z}_{0}\sim p_{data},t\sim\mathcal{U}[k,T],\gamma\sim\mathcal{U}[0,1]}[|\bm{F}_{\bm{\theta}}(\bm{z}_{\gamma,t},c,\gamma,t)-\bm{F}_{{\bm{\theta}}^{-}}(\hat{\bm{z}}^{\Phi}_{1,t-k},c,1,t-k)|],$ (17)
@@
Algorithm 1 Sub-Path Linear Approximation Distillation (SPLAD)
@@
The pseudo-code for SPLAD is presented in Algorithm 1. SPLAD shares a similar training pipeline with consistency models[44, 26] but can be distinguished as it optimizes the sub-path learning objectives based on the SL-ODEs and utilizes the $\gamma$-conditioned training. For a pair of input noise and time step $(\bm{z}_{t},t)$, SPLAM gives the prediction of the denoised latent $\hat{\bm{z}}_{0}$ as:
@@
for one-step generation, adhering strictly to the $\gamma=1$ condition. We also use the same iterative sample strategy as illustrated in [44] which could improve the quality of the generated images. In practice, we set the $-\gamma\text{-}$embedding to $\mathbf{0}$ for $\gamma=1$, thereby allowing the weights associated with trained $\gamma$-embeddings to be discarded post-training. Thus our Sub-Path Linear Approximation Model (SPLAM) necessitates no additional parameters beyond the training phase and can be utilized the same as teacher models.